# EMMA: End-to-End Multimodal Model for Autonomous Driving

### v2

@@ 4 Related Works
End-to-end autonomous driving research enjoys a rich history and has evolved significantly since ALVINN (Pomerleau, 1988) employed shallow neural networks to predict control signals. The field benefited from further deep learning advancements: e.g. DAVE-2 (Bojarski et al., 2016) and ChauffeurNet (Bansal et al., 2019), leveraged deeper neural architectures and incorporated sophisticated perception and motion planning modules respectively. Recent research has expanded to include multimodal inputs (Codevilla et al., 2018; Prakash et al., 2021), multi-task learning (Chitta et al., 2022; Wu et al., 2022), reinforcement learning (Chekroun et al., 2023; Chen et al., 2021; Kendall et al., 2019; Liang et al., 2018; Toromanoff et al., 2020), and distillation (Chen et al., 2020; Zhang and Ohn-Bar, 2021; Zhang et al., 2021). Unified planning frameworks such as VAD (Jiang et al., 2023) and UniAD (Hu et al., 2023) integrated planning with conventional modules in open-loop environments. More studies have been proposed to examine the robustness, safety, and transferability from synthetic environments to real-world domains. However, recent findings from Ego-MLP (Zhai et al., 2023) and BEV-Planner (Li et al., 2024) revealed that these methods could potentially overfit to ego status despite their good performance on benchmark tasks. Our work revisits the simplicity of earlier end-to-end models such as ALVINN and DAVE-2, enhancing them with powerful MLLMs.
@@
Vision language models for autonomous driving have gained increasing interest, focusing on achieving explainable driving behavior and generalizability through end-to-end learning frameworks. DriveGPT4 (Xu et al., 2024) utilizes LLMs to both explain vehicle actions and predict control signals in an iterative Q&A format. Drive Anywhere (Wang et al., 2024b) introduces patch-aligned feature extraction from MLLMs for text-based driving decision queries, while OmniDrive (Wang et al., 2024a) features a 3D vision-language model design for reasoning and planning. Other approaches use MLLMs in graph-based VQA contexts (DriveLM (Sima et al., 2024)), or apply chain-of-thought reasoning (DriveVLM (Tian et al., 2024)) to tackle multiple driving-related tasks. For non-end-to-end models, LLM-Drive (Chen et al., 2024b) leverages LLMs with object-level vector inputs for control planning. In contrast, our work studies the end-to-end fine-tuning of a state-of-the art MLLM for driving tasks, employing a generalist approach that emphasizes open-world driving capabilities.
@@
Multimodal large language models (MLLM) extend LLMs (Vaswani et al., 2017; Devlin, 2019; Raffel et al., 2020; Gemini Team Google, 2023; Reid et al., 2024; Chowdhery et al., 2023; Anil et al., 2023; Radford et al., 2018, 2019; Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023a, b; Dubey et al., 2024) to multiple modalities, leveraging their generalizability, reasoning capabilities, and contextual understanding. Early explorations (Donahue et al., 2015; Vinyals et al., 2015; Chen et al., 2022a) focused on specific vision-language problems or open-set object detection Liu et al. (2024b); Zareian et al. (2021); Gu et al. (2022), while recent research has scaled up both trask diversity and model sizes for improved generalizability and few-shot capabilities (Cho et al., 2021; Chen et al., 2022b; Wang et al., 2022; Lu et al., 2022; Alayrac et al., 2022; Yu et al., 2022; Chen et al., 2023, 2024c; Wang et al., 2024c; Peng et al., 2024; Huang et al., 2023; Lu et al., 2024). Notable examples include Flamingo (Alayrac et al., 2022), a 70B model (Hoffmann et al., 2022) which achieved state-of-the-art quality for multiple few-shot vision benchmarks, and CoCa (Yu et al., 2022) a 2.1B parameter model which demonstrated state-of-the-art performance on zero-shot transfer and various downstream tasks including ImageNet classification. PaLI (Chen et al., 2023, 2024c), at 55B parameters, achieves better performance across multiple vision and language tasks by scaling both the vision and language model components jointly. These early works demonstrate the strong performance and generalizability of MLLMs. Recent trends have seen the integration of native multi-modal inputs in LLMs, such as Gemini (Gemini Team Google, 2023; Reid et al., 2024), GPT-4o, and Llama3-v (Dubey et al., 2024; Liu et al., 2024a). Our work explores the application of these promising new models for generalist end-to-end autonomous driving.
5 Limitations, Risks, and Mitigations
@@
In the previous sections, we demonstrate state-of-the-art end-to-end motion planning on the nuScenes planning benchmark. We also achieve competitive performance for end-to-end motion planning on the WOMD planning benchmark and camera-primary 3D detection on WOD. Furthermore, our generalist setup improved the quality across multiple tasks through joint training. Despite these promising results, we acknowledge the limitations of our work and propose directions for building on this foundation and addressing such challenges in future research.
@@
Verification of the predicted driving signals: Our model can directly predict driving signals without relying on intermediate outputs, such as object detection or road graph estimation. This approach introduces challenges for both real-time and post-hoc verification. We have demonstrated that our generalist model can jointly predict additional human readable outputs such as objects and road graph elements, and the driving decision can be further explained with chain-of-thought driving rationale. However, there is no guarantee that these outputs will be always consistent despite the empirical observations that they are often indeed consistent. Furthermore, additional outputs incur significant runtime latency overhead for deployment.
@@
Sensor simulation for closed-loop evaluation: It is widely recognized that open-loop evaluation may not be strongly correlated with closed-loop performance. To accurately assess an end-to-end autonomous driving system in a closed-loop environment, a comprehensive sensor simulation solution is necessary. However, the computational cost of sensor simulation is often several times higher than that of behavior simulators. This significant cost burden can hinder thorough testing and verification of an end-to-end models unless substantial optimizations are made.
@@
Challenges of onboard deployment: Autonomous driving demands real-time decision-making, which poses a significant challenge when deploying large models due to their increased inference latency. This creates a need for optimizing the model or distilling it into a more compact form suitable for deployment, all while maintaining performance and safety standards. Achieving this delicate balance between model size, efficiency, and quality is crucial for the successful real-world deployment of autonomous driving systems, and represents a key area for future research.
6 Conclusion
In this paper, we present EMMA, a Gemini-powered end-to-end multimodal model for autonomous driving. It treats Gemini as a first class citizen and recasts autonomous driving tasks as vision question answering problems to fit the paradigm of MLLMs, aiming at maximizing the utility of Gemini’s world knowledge and its reasoning capability equipped with chain-of-thought tools. Unlike historical cascaded systems with specialized components, EMMA directly maps raw camera sensor data into various driving-specific outputs, including planning trajectories, perception objects, and road graph elements. All task outputs are represented as plain text and thus can be jointly processed in a unified language space through task-specific prompts. Empirical results show that EMMA achieves state-of-the-art or competitive results on multiple public and internal benchmarks and tasks, including end-to-end planning trajectory prediction, camera-primary 3D object detection, road graph estimation, and scene understanding. We also demonstrate that a single co-trained EMMA can jointly produce the outputs for multiple tasks, while matching or even super-passing the performance of individually trained models, highlighting its potential as a generalist model for many autonomous driving applications.
While EMMA demonstrates promising results, it is still in an early stage with challenges and limitations on onboard deployment, spatial reasoning capability, interpretability, and closed-loop simulation. Despite that, we believe our EMMA findings will inspire further research and progress in the space.
Acknowledgement
We thank many collaborators: Michael Wei, Zoey Yang, Zhenpei Yang, Kan Chen, Ming Ji, Qichi Yang, Yang Wang, Pan Li, Yan Jiao, Haricharan Lakshman, Ruotian Luo, Ningshan Zhang, Anant Subramanian, Chao-Yeh Chen, Xuetao Hao, Andre Cornman, Cheol Park, Xi Chen, and Fei Xia. We also thank the support from Srikanth Thirumalai, Vincent Vanhoucke, Kim Tran, Julia Ilina, Adam Neustein, Ben Loewenstein, Chen Wu, Congcong Li, and Annie Cheng.

### v3

@@ 4 Related Works
End-to-end autonomous driving research enjoys a rich history and has evolved significantly since ALVINN (Pomerleau, 1988) employed shallow neural networks to predict control signals. The field benefited from further deep learning advancements: e.g. DAVE-2 (Bojarski et al., 2016) and ChauffeurNet (Bansal et al., 2019) leveraged deeper neural architectures and incorporated sophisticated perception and motion planning modules respectively. Recent research has expanded to include multimodal inputs (Codevilla et al., 2018; Prakash et al., 2021), multi-task learning (Chitta et al., 2022; Wu et al., 2022), reinforcement learning (Chekroun et al., 2023; Chen et al., 2021; Kendall et al., 2019; Liang et al., 2018; Toromanoff et al., 2020), and distillation (Chen et al., 2020; Zhang & Ohn-Bar, 2021; Zhang et al., 2021). Unified planning frameworks such as VAD (Jiang et al., 2023; Chen et al., 2024c), UniAD (Hu et al., 2023), PARA-Drive (Weng et al., 2024), and GenAD (Zheng et al., 2024) integrated planning with conventional modules in open-loop environments. More studies have been proposed to examine the robustness, safety, and transferability from synthetic environments to real-world domains. However, recent findings from AD-MLP (Zhai et al., 2023) and BEV-Planner (Li et al., 2024) revealed that these methods could potentially overfit to ego status despite their good performance on benchmarks. Our work revisits the simplicity of earlier end-to-end models such as ALVINN and DAVE-2, enhancing them with powerful MLLMs.
@@
Vision language models for autonomous driving have gained increasing interest, focusing on achieving explainable driving behavior and generalizability through end-to-end learning frameworks. DriveGPT4 (Xu et al., 2024) and LMDrive Shao et al. (2024) utilize LLMs to explain vehicle actions and predict control signals in an iterative Q&A format. Drive Anywhere (Wang et al., 2024c) introduces patch-aligned feature extraction from MLLMs for text-based driving decision queries, while OmniDrive (Wang et al., 2024a) features a 3D vision-language model design for reasoning and planning. Other approaches use MLLMs in graph-based VQA contexts ((Sima et al., 2024)), integrate LLMs in a BEV-based planner ((Pan et al., 2024), or apply chain-of-thought reasoning ((Tian et al., 2024; Wang et al., 2024b; Bhattacharyya et al., 2023)) to tackle multiple driving-related tasks. Modular architectures such as LLM-Drive (Chen et al., 2024b) leverage LLMs with object-level vector inputs for planning. In contrast, our work studies end-to-end fine-tuning of a state-of-the art MLLM for driving tasks, employing a generalist approach that emphasizes open-world driving capabilities.
@@
Multimodal large language models (MLLM) extend LLMs (Vaswani et al., 2017; Devlin, 2019; Raffel et al., 2020; Gemini Team Google, 2023; Reid et al., 2024; Chowdhery et al., 2023; Anil et al., 2023; Radford et al., 2018; 2019; Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023a; b; Dubey et al., 2024) to multiple modalities, leveraging their generalizability, reasoning capabilities, and contextual understanding. Early explorations (Donahue et al., 2015; Vinyals et al., 2015; Chen et al., 2022a) focused on specific vision-language problems or open-set object detection Liu et al. (2024b); Zareian et al. (2021); Gu et al. (2022), while recent research has scaled up both trask diversity and model sizes for improved generalizability and few-shot capabilities (Cho et al., 2021; Chen et al., 2022b; Wang et al., 2022; Lu et al., 2022; Alayrac et al., 2022; Yu et al., 2022; Chen et al., 2023; 2024d; Wang et al., 2024d; Peng et al., 2024; Huang et al., 2023; Lu et al., 2024). Notable examples include Flamingo (Alayrac et al., 2022), a 70B model (Hoffmann et al., 2022) which achieved state-of-the-art quality for multiple few-shot vision benchmarks, and CoCa (Yu et al., 2022) a 2.1B parameter model which demonstrated state-of-the-art performance on zero-shot transfer and various downstream tasks including ImageNet classification. PaLI (Chen et al., 2023; 2024d), at 55B parameters, achieves better performance across multiple vision and language tasks by scaling both the vision and language model components jointly. These early works demonstrate the strong performance and generalizability of MLLMs. Recent trends have seen the integration of native multi-modal inputs in LLMs, such as Gemini (Gemini Team Google, 2023; Reid et al., 2024), GPT-4o, and Llama3-v (Dubey et al., 2024; Liu et al., 2024a). Notably, researchers also apply MLLMs to robotic navigation (Zhang et al., 2024; Sun et al., 2024) and manipulation (Brohan et al., 2023; Alakuijala et al., 2024; Wang et al., 2023). Our work explores the application of these promising new models for generalist end-to-end autonomous driving.
5 Conclusion
In this paper, we present EMMA, a Gemini-powered end-to-end multimodal model for autonomous driving. It treats Gemini as a first class citizen and recasts autonomous driving tasks as vision question answering problems to fit the paradigm of MLLMs, aiming at maximizing the utility of Gemini’s world knowledge and its reasoning capability equipped with chain-of-thought tools. Unlike historical cascaded systems with specialized components, EMMA directly maps raw camera sensor data into various driving-specific outputs, including planning trajectories, perception objects, and road graph elements. All task outputs are represented as plain text and thus can be jointly processed in a unified language space through task-specific prompts. Empirical results show that EMMA achieves state-of-the-art or competitive results on multiple public and internal benchmarks and tasks, including end-to-end planning, camera-primary 3D object detection, road graph estimation, and scene understanding. We also demonstrate that a single co-trained EMMA can predict multiple tasks, while matching or even super-passing the performance of individually trained models, highlighting its potential as a generalist model for autonomous driving.
Appendix A Appendix
We organize this appendix section as follows:
  1. 1.
    We summarize all 12 categories in the meta decision in the chain-of-thought reasoning formulation in Table 6.
  2. 2.
    We supply details on the 3D object detection metrics.
  3. 3.
    To provide insights for future development, we present three distinct failure examples in Figure 12.
  4. 4.
    To facilitate reproduction, we provide an example of the concrete prompts used for EMMA Generalist and their corresponding model-predicted answers in Table 7.
  5. 5.
    Despite the promising results, we acknowledge the limitations of our work and propose directions for future research in Section A.5.
A.1 Meta Decision in Chain-of-Thought Reasoning
We describe the chain-of-thought reasoning in Section 2.2 in the main paper. One important component is the meta decision, where we first partition the decision into 12 categories with heuristics and transform them into natural languages.
We summarize the 12 categories in Table 6. We use speed at 3 different future timestamps, i.e., at future 0, 1, and 3 seconds, as the decision points for different categories. If we are able to identify the cause of speed changes, e.g., due to traffic signs or critical objects, then we also append it to the description. We plan to explore more fine-grained meta decisions and reasoning in a future work.
Speed at 0s Speed at 1s Speed at 3s Meta Decision Description
stationary stationary stationary “Stay stationary.”
stationary moving - “Start moving soon.”
stationary stationary moving “Stay stationary for now, then start moving soon.”
moving constant constant “Keep speed.”
moving constant increase “Keep speed, then accelerate.”
moving constant decrease “Keep speed, then brake.”
moving increase increase “Accelerate.”
moving increase constant “Accelerate, then keep high speed.”
moving increase decrease “Accelerate, then brake.”
moving decrease decrease “Brake.”
moving decrease constant “Brake, then keep low speed.”
moving decrease increase “Brake, then accelerate.”
Table 6: Summary of 12 categories of the meta decision in chain-of-thought reasoning. We use speed at 3 different future timestamps, i.e., at future 0, 1, and 3 seconds, as the decision points. If we are able to identify the cause of speed changes, e.g., due to traffic signs or critical objects, then we also append it to the description.
A.2 Distance Breakdowns of 3D Detection Metrics
We plot the distance breakdowns of the camera-primary 3D object detection experiments in Figure 11. We observe that the performance gap between EMMA+ and the baseline models diminishes as the distance to objects increases. We attribute this phenomenon to the potentially lower resolution of the camera input used by EMMA compared to that of the baseline models.
Figure 11: Distance breakdowns of the camera-primary 3D object detection experiments on WOD (Sun et al., 2020) using the standard LET matching (Hung et al., 2024). We observe that the performance gap between EMMA+ and the baseline models diminishes as the distance to objects increases. We attribute this phenomenon to the potentially lower resolution of the camera input used by EMMA compared to that of the baseline models.
A.3 Failure Examples
While Section 3.6 showcases numerous successful predictions from EMMA across trajectory, detection, and road graph tasks, it is equally important to analyze its limitations. To provide insights for future development, we present three distinct failure scenarios in Figure 12, each illustrating a different type of error.
Refer to captionRefer to captionRefer to caption
EMMA’s predicted trajectory exhibits a suboptimal clearance from the motorbike. While the prediction is validated when the motorbike later tracks further right, the ideal path would have included a greater initial lateral offset.
Refer to captionRefer to captionRefer to caption
EMMA fails to detect a distant oncoming vehicle, though it is correctly identified in the subsequent frame. This one-frame detection delay is suboptimal, as the ego vehicle plans to nudging to the left. Earlier detection would have led to a safer and more controlled trajectory.
Refer to captionRefer to captionRefer to caption
While EMMA correctly identifies the immediate drivable lanes, it fails to detect an upcoming lane merge further down the road. This oversight may be attributed to the degraded painted arrow.
Figure 12: Failure examples of EMMA prediction. Each row contains a scenario where EMMA’s predictions were partially flawed: end-to-end planning trajectory (left), 3D object detection (middle), and road graph estimation (right).
A.4 Concrete Prompts and Answers
As demonstrated in Section 2.3 and Figure 2, EMMA is capable of generating desired outputs based on varying prompts. To facilitate reproducibility, Table 7 provides the specific prompts used and their corresponding model-predicted answers.
Image Prompt and Answer
[Uncaptioned image] Prompt: Assume I am driving at the coordinate 0.0, 0.0. The high-level behavior attention is: go straight. The past trajectory under vehicle coordinate is: -4.09, 0.01 and -3.94, 0.01 and -3.71, 0.01 and -3.42, 0.01 and -3.05, 0.00 and -2.60, 0.00 and -2.07, -0.00 and -1.45, -0.00 and -0.76, -0.00 and 0.00, 0.00. The past ego velocity under vehicle coordinate is: 0.94, -0.01 and 1.29, -0.01 and 1.66, -0.01 and 2.04, -0.02 and 2.45, -0.02 and 2.88, -0.02 and 3.26, -0.02 and 3.65, -0.01 and 3.98, -0.00. The past ego acceleration under vehicle coordinate is: 0.32, -0.00 and 0.34, -0.00 and 0.37, -0.00 and 0.38, -0.00 and 0.40, -0.00 and 0.43, 0.00 and 0.38, 0.00 and 0.39, 0.01 and 0.34, 0.01. What is my future trajectory in next 5 seconds under vehicle coordinate?
Answer: 0.83, 0.01 and 1.72, 0.01 and 2.67, 0.02 and 3.67, 0.04 and 4.73, 0.05 and 5.84, 0.07 and 6.95, 0.09 and 8.08, 0.11 and 9.23, 0.13 and 10.41, 0.15 and 11.62, 0.17 and 12.87, 0.19 and 14.15, 0.22 and 15.47, 0.24 and 16.80, 0.26 and 18.15, 0.27 and 19.52, 0.28 and 20.92, 0.28 and 22.36, 0.28 and 23.83, 0.28 and 25.34, 0.27 and 26.89, 0.26 and 28.49, 0.25 and 30.14, 0.23 and 31.83, 0.22
[Uncaptioned image] Prompt: What are the 3D bounding boxes for objects in the following list: vehicle, pedestrian, and cyclist? Describe the bounding boxes in the format of [center_x, center_y, bottom_z, width, length, height, heading, class], where heading is in radian, class in text, and other dimensions are in meter?
Answer: -12.91 -9.23 -0.21 12.99 3.21 3.45 -2.25 vehicle and 21.81 -4.15 -0.15 4.61 2.10 1.65 -2.18 vehicle and 25.61 -1.11 -0.12 1.89 0.86 1.86 -0.23 cyclist and 39.12 3.25 -0.13 4.39 2.13 1.79 3.14 vehicle and 45.51 3.18 -0.12 4.65 2.10 1.54 3.14 vehicle and 43.81 21.51 -0.01 4.71 2.19 1.51 -2.16 vehicle and 40.01 20.11 0.09 4.92 2.39 2.02 -2.17 vehicle and 44.81 3.16 -0.11 4.51 2.10 1.55 3.14 vehicle and 49.11 13.39 0.02 1.11 0.83 1.68 -0.05 pedestrian and 52.31 3.02 0.05 4.85 2.23 1.92 -3.14 vehicle and 51.11 15.41 0.05 1.12 0.85 1.78 -0.01 pedestrian and 62.51 3.35 0.49 4.51 2.15 1.85 -3.14 vehicle and 72.11 2.99 0.92 4.45 2.13 1.74 -3.13 vehicle and 79.21 3.41 1.19 4.41 2.14 1.72 -3.13 vehicle and 85.51 3.35 1.45 4.36 2.08 1.60 -3.13 vehicle and 85.11 11.11 1.41 4.49 2.15 1.76 3.11 vehicle and 92.21 3.51 1.69 4.45 2.18 1.90 -3.13 vehicle
[Uncaptioned image] Prompt: My current position is 1.28, -0.00. Where are driveable lanes in front of me? Describe the lanes as a series of points relative to 0.00, 0.00?
Answer: (0.00, -0.11 and 10.00, -0.11 and 20.00, -0.11 and 30.00, -0.11 and 40.00, -0.11 and 51.11, -0.11 and invalid and invalid and invalid and invalid) valid ; (51.11, -0.11 and 61.10, -0.06 and 71.09, -0.01 and 81.08, 0.04 and 91.07, 0.09 and invalid and invalid and invalid and invalid and invalid) valid ; (0.00, -0.12 and 10.00, -0.19 and 15.11, -0.22 and 21.99, 0.62 and 27.72, 2.16 and 33.11, 4.65 and 38.02, 8.00 and 42.35, 12.05 and invalid and invalid) valid ; (0.00, -0.12 and 10.00, -0.13 and 17.11, -0.14 and 23.01, 0.58 and 28.76, 2.09 and 34.16, 4.56 and 39.07, 7.91 and 43.43, 11.95 and invalid and invalid) valid ; (0.00, -0.12 and 10.00, -0.13 and 16.11, -0.14 and 22.01, 0.55 and 27.76, 2.05 and 33.14, 4.55 and 38.00, 7.98 and 42.29, 12.09 and invalid and invalid) valid ; (0.00, -3.31 and 10.00, -3.25 and 20.00, -3.19 and 30.00, -3.13 and 40.00, -3.07 and 49.35, -3.01 and invalid and invalid and invalid and invalid) valid ; (invalid and invalid and invalid and invalid and invalid and invalid and invalid and invalid and invalid and invalid) invalid …
Table 7: The specific prompts used and their corresponding model-predicted answers. Numerical values are color-coded in blue and predicted separators in red for better visualization.
A.5 Limitations, Risks, and Mitigations
@@
In the main paper, we demonstrate state-of-the-art end-to-end motion planning on the nuScenes planning benchmark. We also achieve competitive performance for camera-primary 3D detection on WOD. Furthermore, our generalist setup improves the quality across multiple tasks through joint training. Despite these promising results, we acknowledge the limitations of our work and propose directions for building on this foundation and addressing such challenges in future research.
@@
Verification of the predicted driving signals: Our model can directly predict driving signals without relying on intermediate outputs, such as object detection or road graph estimation. This approach introduces challenges for both real-time and post-hoc verification. We have demonstrated that our generalist model can jointly predict additional human readable outputs such as objects and road graph elements, and the driving decision can be further explained with chain-of-thought driving rationale. However, there is no guarantee that these outputs will be always consistent despite the empirical observations that they are often indeed consistent. Improving driving rationale is one of our future research directions.
@@
Sensor simulation for closed-loop evaluation: To accurately assess an end-to-end autonomous driving system in a closed-loop environment, a comprehensive sensor simulation solution is necessary. However, the computational cost of sensor simulation is often much higher than that of behavior simulators. This significant cost burden can hinder thorough testing and verification of an end-to-end models; however, the field of efficient sensor simulation methods continues to rapidly evolve, with the promise of significantly mitigating this burden.
More targeted and accurate open-loop evaluation: While existing open-loop evaluation offers low computational cost, its results are sometimes unreliable. The popular nuScenes (Caesar et al., 2020) benchmark, for instance, exhibits well-documented limitations. Key metrics like collision rate are sensitive to hyperparameter choices, such as the BEV grid resolution (Weng et al., 2024; Zhai et al., 2023). Furthermore, many of its scenarios lack planning diversity and can be trivially solved by extrapolating historical trajectories. Initiatives like NAVSIM (Dauner et al., 2024) are beginning to address these vulnerabilities, but this highlights a critical need for future work on developing more challenging and trustworthy open-loop evaluation frameworks.
@@
Challenges of onboard deployment: Autonomous driving demands real-time decision-making, which poses a significant challenge when deploying large models due to their increased inference latency. This creates a need for optimizing the model or distilling it into a more compact form suitable for deployment, all while maintaining performance and safety standards.
While the main paper is focused on establishing the EMMA architecture, the framework is amenable to various well-established optimization techniques. To illustrate this adaptability, we have experimented with a latency-optimized EMMA configuration. By incorporating strategies such as SARA-RT (Leal et al., 2024), employing shorter action sequences, streamlining by removing explicit reasoning chains, etc., this variant achieves an inference speed of 3 FPS, a 67% speedup compared to UniAD’s 1.8 FPS. This preliminary result underscores that the core EMMA architecture possesses the flexibility to be effectively adapted into specialized, low-latency variants. Achieving a delicate balance between model size, efficiency, and quality is crucial for the successful real-world deployment of autonomous driving systems, and represents a key area for future research.